# Haitao Xu    Jingru Zhang

### v1

@@ 1 Introduction
Facility location problems with uncertain demands have been explored a lot [7, 10, 13, 14, 1, 2] due to the inherent uncertainty of collected demand data. In this paper, we study a facility location problem with uncertain demands, that is, the two-center problem on a cactus graph network w.r.t. a set of uncertain points each with $O(m)$ possible locations on the network.
On tree networks, this problem was solved in $O(mn\log mn)$ time by a prune-and-search algorithm [17]. On cactus graphs, if $m=1$, the problem was addressed in $O(n\log^{2}n)$ time [4]; for $m>1$, as far as we are aware, only the one-center problem has been explored [9].
Let $G=(V,E)$ be the given (undirected) cactus graph. For any two vertices $u,v$, denote by $e(u,v)$ their edge. Any point $q\in e(u,v)$ is described by a tuple $(u,v,t)$ where $t$ is the distance of $q$ to $u$ along $e(u,v)$. For any two points $p$ and $q$ on $G$, their distance $d(p,q)$ is exactly their shortest path length.
@@
For any point $q$ on $G$, the distance between any $P_{i}\in\mathcal{P}$ and $q$ is the expected version, say $\mathsf{E}\mathrm{d}(P_{i},q)$, and it is defined as $\sum_{j=1}^{m}f_{ij}\cdot d(p_{ij},q)$. For any two points $q_{1},q_{2}$ on $G$, we define $\phi(q_{1},q_{2})=\max_{P_{i}\in\mathcal{P}}w_{i}\min\{\mathsf{E}\mathrm{d}(P_{i},q_{1}),\mathsf{E}\mathrm{d}(P_{i},q_{2})\}$. The problem aims to compute two points on $G$, i.e., centers $q^{*}_{1}$ and $q^{*}_{2}$, so as to minimize $\phi(q_{1},q_{2})$.
@@
Another most related problem is the general $k$-center problem where $k\geq 2$ centers are placed on the network so as to minimize the maximum (weighted) distance between demands and centers. For $m=1$, the $k$-center problem on trees was recently addressed in $O(n\log n)$ time by Wang and Zhang [15]. Bai et al. [3] gave an $O(n^{2}k^{2})$-time algorithm for the problem on cactus graphs. When $m>1$, Wang and Zhang [14] proposed an $O(n^{2}\log n\log mn+mn\log^{2}mn\log n)$ time algorithm for this problem on tree networks. But no previous work exists for the cactus version except for $k=1$ [9].
@@
Our algorithm is a generalized version of the decision algorithm [17] for $G$ being a tree. Specifically, we first transform $G$ into its tree representation $T$ where each node uniquely represents a cycle, a hinge (that is a vertex on a cycle of degree more than $2$), or an out-of-cycle vertex. Then, at most two ‘binary searches’ are performed on $T$ where each finds a distinct out-of-cycle edge or a cycle on $G$ that must contain a peripheral center. Last, we decide whether two centers can be placed on the obtained peripheral-center edges or cycles to cover $\mathcal{P}$ under $\lambda$. The running time is dominated by the $O(m^{2}n^{2})$ time to handle the case where placing two centers on a cycle to cover $\mathcal{P}$ under $\lambda$.
@@ 2 Preliminaries
Lemma $5$ in [9] reveals that any general instance of the one-center problem w.r.t. $\mathcal{P}$ on $G$ can be reduced into a vertex-constrained instance. For the two-center problem, we have a similar result.
@@
The reduction is similar to Lemma $5$ in [9]. For completeness and correctness, we include the details in the following. First, join into $G$ a vertex for every point interior of an edge that holds locations, and assign all its locations to this vertex. Second, repeat to remove from $G$ every empty cycle except for its only hinge as well as empty vertices of degree $1$. Then, for each cycle whose only hinge is non-empty, remove its other vertices except for its hinge. Clearly, these removed components of $G$ contain no centers.
Further, for each cycle with only two hinges, if all its other vertices are empty, replace this cycle by an edge incident to its two hinges of length equal to their shortest path along this cycle. Because it is proved in Lemma $5$ [9] that the center of $\mathcal{P}$ on $G$ is not interior of their longest path on the cycle. Clearly, it also applies to our problem since the center that leads $\lambda^{*}$ is the center of a subset of uncertain points of $\mathcal{P}$. Last, keep removing every empty vertex of degree $2$ by connecting its two adjacent vertices directly.
At this point, the size of $G$ is $O(mn)$ and so it contains $O(mn)$ empty vertices (each of degree at least $3$). Create $O(m)$ additional locations of zero probabilities for each $P_{i}\in\mathcal{P}$. It follows that each $P_{i}\in\mathcal{P}$ has $O(m)$ locations and every location is at a vertex of $G$ and every vertex has at least one location. The reduction takes $O(|G|+mn)$ time. It is clear to see that solving the two-center problem w.r.t. $\mathcal{P}$ on $G$ is equivalent to solving the obtained vertex-constrained instance. ∎
@@
A vertex in a cycle of degree at least $3$ is called a hinge, and an out-of-cycle vertex is called a $G$-vertex. Any cactus graph can be represented by a tree that is generated by replacing every cycle with a node and then connecting it with all its hinges by edges of length zero (see Fig. 1). Such a tree representation can be built in linear time [6, 9].
@@
Figure 1: (a) Illustrating a cactus graph that consists of three cycles $\{C_{1},C_{2},C_{3}\}$, two hinges $\{h_{1},h_{2}\}$ (squares), and six G-vertices (disks). Cycle $C_{2}$ has four split subgraphs $\{G_{1},G_{2},G_{3},G_{4}\}$, and four hanging subgraphs $\{C_{1},C_{2},G_{3}\cup\{h_{2}\},G_{4}\cup\{h_{2}\}\}$. (b) Illustrating the tree representation of the cactus graph in (a) where each cycle is uniquely represented by a (circular) node. Specifically, node $u_{1}$ is for cycle $C_{1}$, node $u_{2}$ for cycle $C_{2}$, and node $u_{3}$ for cycle $C_{3}$. Additionally, $u_{2}$ and its all adjacent hinge nodes induce its $H$-subtree $T^{H}_{u_{2}}$. Every hanging subgraph of $C_{2}$ is uniquely represented by a hanging subtree of $T^{H}_{u_{2}}$.
@@
For any $P_{i}\in\mathcal{P}$, the sum of probabilities of all $P_{i}$’s positions in a subgraph $G^{\prime}$ of $G$ is referred to as the probability sum of $P_{i}$ in $G^{\prime}$. The median of $P_{i}$, denote by $x_{i}^{*}$, is the point on $G$ that minimizes $\mathsf{E}\mathrm{d}(P_{i},x)$. The following lemmas are proved in the one-center algorithm [9].
@@
    Given any point $x$ on $G$, computing for each $1\leq i\leq n$ the expected distance $\mathsf{E}\mathrm{d}(P_{i},x)$ in $O(mn)$ time;
@@ 3 The Decision Algorithm
Given any value $\lambda>0$, the decision problem is to determine whether two centers can be placed on $G$ so that the objective value is no more than $\lambda$. If yes, $\lambda\geq\lambda^{*}$ and it is feasible; otherwise, $\lambda<\lambda^{*}$ so it is infeasible.
For any point $x$ of $G$, we say that $x$ can cover $\mathcal{P}$ (under $\lambda$) if $\phi(x,x)\leq\lambda$. An out-of-cycle edge of $G$ is called a peripheral-center edge if a center can be placed on this edge to cover all uncertain points whose medians are in a subgraph generated by removing it except for its incident vertices from $G$, and they cannot be covered by any point not in this subgraph and the edge. Clearly, placing a center in a peripheral-center edge leads least centers needed to cover $\mathcal{P}$.
@@
The decision algorithm [16] for $G$ be a tree performs ‘binary searches’ on $G$’s tree representation to find two peripheral-center edges. It then determines whether two centers can be placed on the obtained peripheral-center edge(s) to cover $\mathcal{P}$.
However, a cactus graph may have cycles, so we must consider the case where centers must be placed in cycles. We say that a cycle of $G$ is a peripheral-center cycle if all uncertain points whose medians are in all but one hanging subgraphs of the cycle can be covered by some points of the cycle but some of them cannot be covered by any point of that excluded hanging subgraph except for its hinge in the cycle.
@@
Our algorithm generalizes the decision algorithm [16] for trees: First, it finds two minimum subtrees on $T$ that respectively contain a peripheral-center edge or cycle on $G$. Next, we determine if two centers can be placed on their represented subgraph(s) of $G$ to cover $\mathcal{P}$ under $\lambda$. By the definitions, a peripheral-center edge or cycle must contain centers in order to place least centers on $G$ to cover $\mathcal{P}$. It follows that if two points of the peripheral-center edges and cycles can cover $\mathcal{P}$ then $\lambda\geq\lambda^{*}$ and otherwise, $\lambda<\lambda^{*}$.
@@
Let $c^{1}$ and $c^{2}$ be any two peripheral-center edges or cycles on $G$. Note that $c^{1}=c^{2}$ if all necessary centers are on the same out-of-cycle edge or cycle. Use $T^{1}$ and $T^{2}$ to represent the two subtrees of $T$ that respectively consist of $c^{1}$ and $c^{2}$. At the beginning, we have $T^{1}=T^{2}=T$. Also, assign a flag $f$ to every node $u$ of $T^{1}$ and $T^{2}$ so that $u$’s $f$ is true if and only if the hanging subtrees of $u$ in $T/T^{1}$ contain peripheral-center edges or cycles. The flag $f$ of every node in $T^{1}$ is initialized as false.
@@
We shall now present how to find a peripheral-center edge or cycle node on $T^{1}$ and then $T^{2}$ in turn. The procedure needs the supports of Lemmas 5 and 6 whose proofs are delayed to Section 3.1.
@@
Compute the centroid $c$ on $T^{1}$ which is a node of $T^{1}$ whose hanging subtrees each is of size no more than half of $T^{1}$ and can be found in $O(|T^{1}|)$ time [11, 12]. Then, if $c$ is a hinge or $G$-node, by applying Lemma 5 to $c$, we determine whether a center must be placed at $G_{c}$ and if not, which hanging subtree of $c$ contains a peripheral-center edge or cycle. If only $c$ is obtained, which means that $G_{c}$ must be a center, then we set $T^{1}$ as $c$.
@@
Generally, Lemma 5 returns one or two nodes that are adjacent to $c$ in $T^{1}$. In the former case where only one adjacent node of $c$ is returned, the corresponding hanging subtree of $c$ on $T$ contains all necessary centers. Hence, we find in $T^{1}$ $c$’s hanging subtree including this node and reset $T_{1}$ as this subtree. The latter case means that the two corresponding hanging subtrees of $c$ on $T$ each contains a peripheral-center edge or cycle. Find in $O(|T^{1}|)$ time the two hanging subtrees of $c$ in $T^{1}$ respectively containing one of the two nodes. Let $T^{1}$ be the one of them that excludes any node with a true flag. Last we set $c$’s flag $f$ in $T^{1}$ as true.
@@
On the other hand, $c$ is a cycle node. Instead, we apply Lemma 6 to decide whether $G_{c}$ is a peripheral-center cycle and which hanging subtree of $c$’s $H$-subtree $T^{H}_{c}$ in $T^{1}$ contains a peripheral-center edge or cycle. As described in the proof for Lemma 6, if $c$ is obtained, which means that $G_{c}$ is a peripheral-center cycle, then we let $T^{1}$ be $c$. If one of its adjacent hinge nodes is returned, which can be known in $O(|T^{1}|)$ time, then a center must be placed at that hinge; hence we set $T^{1}$ as this hinge node.
@@
Otherwise, we obtain two nodes that are respectively in two split subtrees of $T^{H}_{c}$ and adjacent to $c$’s hinge node(s), or only one such node. For the former situation, we find in linear time a hanging subtree of $T^{H}_{c}$ in $T^{1}$ that includes one of the two nodes but excludes any node with a true flag $f$; next we let $T^{1}$ be this subtree and set the flag $f$ as true of $c$’s hinge node in $T^{1}$. For the last case, we set $T^{1}$ as the hanging subtree of $T^{H}_{c}$ in $T^{1}$ that contains this node.
@@
We recursively perform the above procedure to find a peripheral-center edge or cycle on $T^{1}$ until $T^{1}$ consists of only one node or one edge. Clearly, it takes $O(\log mn)$ recursive steps, so the running time is $O(mn\log mn)$.
@@
    $T^{1}$ is a hinge node or $G$-node. A center must be placed at vertex $G_{T^{1}}$. We can decide the feasibility of $\lambda$ in $O(mn\log mn)$ time as follows. Set the weight of every uncertain point that can be covered by $G_{T^{1}}$ as zero. Then we apply the one-center algorithm [9] to solve the one-center problem w.r.t. $\mathcal{P}$ on $G$ in $O(mn\log mn)$ time. If the obtained objective value is not greater than $\lambda$, then $G_{T^{1}}$ and the obtained center can cover $\mathcal{P}$. Hence, $\lambda$ is feasible. Otherwise, $\lambda$ is not feasible.
@@
    $T^{1}$ is a cycle node. By the definition, we first determine which hanging subtree of $T^{1}$’s $H$-subtree contains a peripheral-center edge or cycle. If none, then all necessary centers can be placed on $G_{T^{1}}$. Hence, Lemma 8 presented in Section 3.1 is called to decide the feasibility of $\lambda$ in $O(m^{2}n^{2})$ time. Otherwise, we let $T^{2}$ be that hanging subtree of $T^{1}$’s $H$-subtree in $T$ and set the flag $f$ as true for the hinge node in $T^{2}$ that is adjacent to $T^{1}$ in $T$, which can be carried out in $O(|T|)$ time.
@@
    $T^{1}$ is an edge. If it is between a hinge and a cycle node, clearly, this cycle node represents a peripheral-center cycle. We set $T^{1}$ as this cycle node and then handle this case in the above way in $O(m^{2}n^{2})$ time. Otherwise, $T^{1}$ is an out-of-cycle edge on $G$. We call Lemma 7 in Section 3.1 to decide the feasibility of $\lambda$ in $O(mn\log mn)$ time.
In general, $T^{1}$ is a cycle node representing a peripheral-center cycle, i.e., $c^{1}$, and $T^{2}$ is a subtree of $T$ excluding $T^{1}$ where its one leaf connects it and $T^{1}$ on $T$ and has flag $f$ being true. $G_{T^{2}}$ contains all other necessary centers under $\lambda$.
Proceed with finding a peripheral-center edge or cycle on $T^{2}$ recursively as the above way. After $O(\log mn)$ recursive steps, $T^{2}$ shrinks down to a node or an edge on $T$. If it is an out-of-cycle edge, we call Lemma 7 to decide the feasibility of $\lambda$ in $O(mn\log mn)$ time. Otherwise, $T^{2}$ contains exactly one cycle node, which is not same as $T^{1}$. Let $T^{2}$ be the cycle node, which is $c^{2}$. Subsequently, Lemma 9 in Section 3.1 is applied to $T^{1}$ and $T^{2}$, which decides in $O(mn^{2})$ time the feasibility of $\lambda$.

### v2

@@ 1 Introduction
Facility location problems with uncertain demands have been explored a lot [7, 10, 15, 16, 1, 2] due to the inherent uncertainty of collected demand data. In this paper, we study a facility location problem with uncertain demands, that is, the two-center problem on a cactus graph network w.r.t. a set of $n$ uncertain points each of which has $O(m)$ possible locations on the network.
Let $G=(V,E)$ be the given (undirected) cactus graph, which is an undirected graph where no cycles share edges. For any two vertices $u,v$, denote by $e(u,v)$ their edge and let $l(e(u,v))$ be its length. Any point $q\in e(u,v)$ is described by a tuple $(u,v,t)$ where $t$ is the distance of $q$ to $u$ along $e(u,v)$. Moreover, for any two points $p$ and $q$ on $G$, their distance $d(p,q)$ is the length of their shortest path.
@@
For any point $q$ on $G$, the distance between any $P_{i}\in\mathcal{P}$ and $q$ is the expected version, denoted by $\mathsf{E}\mathrm{d}(P_{i},q)$, and it is defined as $\sum_{j=1}^{m}f_{ij}\cdot d(p_{ij},q)$. For any two points $q_{1},q_{2}$ on $G$, we define $\phi(q_{1},q_{2})=\max_{P_{i}\in\mathcal{P}}w_{i}\cdot\min\{\mathsf{E}\mathrm{d}(P_{i},q_{1}),\mathsf{E}\mathrm{d}(P_{i},q_{2})\}$. The problem aims to compute two points on $G$, that is, centers $q^{*}_{1}$ and $q^{*}_{2}$, so as to minimize $\phi(q_{1},q_{2})$.
When $G$ is a tree network, this problem can be solved in $O(mn\log mn)$ time by the prune-and-search algorithm [19]. On cactus graphs, if $m=1$, the problem was addressed in $O(n\log^{2}n)$ time [4]. But for $m>1$, as far as we are aware, only the one-center problem has been explored [9]. In this paper, we give the first algorithm that solves the two-center problem of uncertain points on a cactus graph in $O(|G|+m^{2}n^{2}\log mn)$ time where $|G|$ is the size of $G$, i.e., $|G|=O(|V|+|E|)$.
@@
Another of most related problems is the general $k$-center problem where $k\geq 2$ centers are placed on the network so as to minimize the maximum (weighted) expected distance between uncertain points and centers. For $m=1$, the $k$-center problem on trees was recently addressed in $O(n\log n)$ time by Wang and Zhang [17]. Bai et al. [3] gave an $O(n^{2}k^{2})$-time algorithm for the problem on cactus graphs. When $m>1$, Wang and Zhang [16] proposed an $O(n^{2}\log n\log mn+mn\log^{2}mn\log n)$ time algorithm for this problem on tree networks. But no previous work exists for the cactus version except for $k=1$ [9].
In the plane, when $m=1$, the two-center problem under the Euclidean metric can be solved in $O(n\log^{2}n)$ time by Wang [14]. For $m>1$, the general $k$-center problem on uncertain points is NP-hard due to the NP-hardness of the case $m=1$ [13] and approximation algorithms have been investigated in [10, 2].
@@
Our algorithm is a generalized version of the decision algorithm [19] for $G$ being a tree. Specifically, we first transform $G$ into its tree representation $T$ where each node uniquely represents a cycle, a hinge (that is a vertex on a cycle of degree more than $2$), or a $G$-vertex (that is a vertex out of any cycle). Then, at most two ‘binary searches’ are performed on $T$ where each finds a distinct out-of-cycle edge or a cycle on $G$ that must contain a peripheral center. Last, we decide whether two centers can be placed on the obtained peripheral-center edges or cycles to cover $\mathcal{P}$ under $\lambda$. The running time is dominated by the $O(m^{2}n^{2})$ time to handle the case where placing two centers on a cycle to cover $\mathcal{P}$ under $\lambda$.
@@ 2 Preliminaries
Lemma $5$ in [9] reveals that any general instance of the one-center problem w.r.t. $\mathcal{P}$ on $G$ can be reduced into a vertex-constrained instance. For the two-center problem, we have the similar result.
@@
The reduction is similar to Lemma $5$ in [9]. For completeness and correctness, we include the details in the following. First, we join into $G$ a vertex for every point that holds locations but is not at any vertex, and assign all locations at the point to the new vertex created for it.
We say that a subgraph of $G$ is empty if it is free of any locations. Next, starting from any out-of-cycle vertex or any hinge, we perform a depth-first search on $G$ to shrink $G$ as follows. For every cycle of only one hinge, if its every non-hinge vertex is empty, we remove this cycle from $G$ except for its only hinge. Also, we remove from $G$ every empty vertex of degree $1$ including the generated by removing such above cycles and empty vertices. Clearly, these removed components of $G$ contain no centers.
Furthermore, for each cycle with only two hinges, if its non-hinge vertices are all empty, we remove the cycle from $G$ by connecting its two hinges directly with an edge of length equal to their shortest-path length along this cycle and then pruning all its non-hinge vertices. Because it was proved in Lemma $5$ [9] that the center of $\mathcal{P}$ on $G$ is not interior of the longer path between two hinges in such a cycle. Clearly, this result also applies to our problem since the center that leads $\lambda^{*}$ is the center w.r.t. a subset of uncertain points in $\mathcal{P}$.
Last, we perform a depth-first search on $G$ as the above to remove every empty vertex of degree $2$ by connecting its two adjacent vertices directly.
At this point, the size of $G$ is $O(mn)$ and so it contains $O(mn)$ empty vertices (each of degree at least $3$). Create $O(m)$ additional locations of zero probabilities for each $P_{i}\in\mathcal{P}$ and assign each of them to a unique empty vertices of $G$. It follows that each $P_{i}\in\mathcal{P}$ has $O(m)$ locations and every location is at a vertex of $G$ and every vertex has at least one location. The reduction takes $O(|G|+mn)$ time. It is clear to see that solving the two-center problem w.r.t. $\mathcal{P}$ on $G$ is equivalent to solving the obtained vertex-constrained instance. ∎
@@
A vertex in a cycle of degree at least $3$ is called a hinge, so any cactus graph can be decomposed at its hinges into cycles and tree subgraphs called grafts. A $G$-vertex is a vertex (in a graft subgraph) that is not included in any cycle. Any cactus graph can be represented by a tree that is generated by replacing every cycle with a node and then connecting it with all its hinges by edges of length zero (see Fig. 1). Such a tree representation can be built in linear time [6, 9].
@@
Figure 1: (a) Illustrating a cactus graph that consists of three cycles $\{C_{1},C_{2},C_{3}\}$, two hinges $\{h_{1},h_{2}\}$ (squares), and six G-vertices (disks). Cycle $C_{2}$ has four split subgraphs $\{G_{1},G_{2},G_{3},G_{4}\}$, and four hanging subgraphs $\{C_{1},C_{3},G_{3}\cup\{h_{2}\},G_{4}\cup\{h_{2}\}\}$. (b) Illustrating the tree representation of the cactus graph in (a) where each cycle is uniquely represented by a (circular) node. Specifically, node $u_{1}$ is for cycle $C_{1}$, node $u_{2}$ for cycle $C_{2}$, and node $u_{3}$ for cycle $C_{3}$. Additionally, $u_{2}$ and its all adjacent hinge nodes induce its $H$-subtree $T^{H}_{u_{2}}$. Every hanging subgraph of $C_{2}$ is uniquely represented by a hanging subtree of $T^{H}_{u_{2}}$.
@@
For any $P_{i}\in\mathcal{P}$, the sum of probabilities of all $P_{i}$’s positions in a subgraph $G^{\prime}$ of $G$ is referred to as the probability sum of $P_{i}$ in $G^{\prime}$. The median of $P_{i}$ is a point on $G$ that minimizes $\mathsf{E}\mathrm{d}(P_{i},x)$. Note that the median of $P_{i}$ may not be unique but let $x_{i}^{*}$ denote any of them. The following lemmas are proved in the one-center algorithm [9].
@@
    Given any point $x$ on $G$, computing for all $1\leq i\leq n$ the expected distance $\mathsf{E}\mathrm{d}(P_{i},x)$ in $O(mn)$ time;
@@ 3 The Decision Algorithm
Given any value $\lambda>0$, the decision problem is to determine whether two centers can be placed on $G$ so that the objective value is no more than $\lambda$. If yes, $\lambda\geq\lambda^{*}$ and it is feasible; otherwise, $\lambda<\lambda^{*}$ so it is infeasible. For any point $x$ of $G$, we say that $x$ covers a subset $\mathcal{P}^{\prime}$ of $\mathcal{P}$ (under $\lambda$) iff $\max_{P_{i}\in\mathcal{P}^{\prime}}w_{i}\cdot Ed(P_{i},x)\leq\lambda$. In other words, the decision problem aims to find at most two points (to place centers) on $G$ that cover $\mathcal{P}$ completely.
We say that an edge $e$ of $G$ is a peripheral-center edge (w.r.t. $\lambda$) if it satisfies the following conditions: (1) $e$ is not in any cycle; (2) $e$ has a hanging subgraph, say $G^{\prime}$, so that the subset of all uncertain points that have their probability sums in $G^{\prime}$ larger than $0.5$ can be covered by a point of $e$; (3) an uncertain point whose median is in $G^{\prime}$ cannot be covered by any point in $G/\{G^{\prime}\cup e\}$. Lemma 2 implies that for each $P_{i}$ whose median is in $G^{\prime}$, $\mathsf{E}\mathrm{d}(P_{i},x)$ is monotonically increasing as $x$ moves along $e$ away from $G^{\prime}$ and $\mathsf{E}\mathrm{d}(P_{i},x)$ at any point of $G/\{G^{\prime}\cup e\}$ is not less than its value at $e$’s incident vertex not in $G^{\prime}$. Hence, in order to place least centers to cover $\mathcal{P}$, a center must be placed on such a peripheral-center edge $e$ to cover all uncertain points with medians in $G^{\prime}$ and most uncertain points with medians in $G/G^{\prime}$; such a center on it is called a peripheral center (in that all other necessary centers are placed in its hanging subgraph excluding $G^{\prime}$).
@@
Clearly, peripheral-center edges are “outermost” subgraphs of smallest size on $G$ that must contain (peripheral) centers under $\lambda$. Hence, the decision algorithm [18] for $G$ being a tree performs “binary searches” on $G$ to find two peripheral-center edges, and then determines whether two points can be found on the peripheral-center edges to cover $\mathcal{P}$; if yes, $\lambda$ is feasible and otherwise, more than two centers are needed so $\lambda$ is infeasible.
In general, $G$ contains cycles. Besides of peripheral-center edges, $G$ might have “outermost” cycles that must contain centers under $\lambda$. Similarly, a cycle $C$ of $G$ is a peripheral-center cycle under $\lambda$ if it satisfies the following conditions: (1) $C$ has at most one hanging subgraph, say $G^{\prime\prime}$, so that the subset of uncertain points that have probability sums in $G^{\prime\prime}$ larger than $0.5$ cannot be covered by a point of $C$; (2) an uncertain point with its median in $C\cup\{G/G^{\prime\prime}\}$ cannot be covered by any point of $G^{\prime\prime}/\{G^{\prime\prime}\cap C\}$, where $G^{\prime\prime}\cap C$ is their common vertex. Because each uncertain point with its median in $C\cup\{G/G^{\prime\prime}\}$ has its expected distance at any point of $G^{\prime\prime}/\{G^{\prime\prime}\cap C\}$ not less than its expected distance at vertex $G^{\prime\prime}\cap C$. In order to place least centers covering $\mathcal{P}$, one must place center(s) in such a peripheral-center cycle $C$ to cover the subset of uncertain points satisfying condition (2) and most uncertain points not in this subset, and other necessary centers must be placed in $C\cup G^{\prime\prime}$.
@@
Our algorithm is a generalization of the decision algorithm [18] for trees: First, it finds on the tree representation $T$ of $G$ two smallest subtrees that each represents a peripheral-center edge or cycle on $G$ under $\lambda$. Next, we determine if two centers can be placed on the obtained peripheral-center edge(s) and cycle(s) to cover $\mathcal{P}$ under $\lambda$. If there exist two points on them that can cover $\mathcal{P}$, then $\lambda\geq\lambda^{*}$, and otherwise, more than two points are needed to cover $\mathcal{P}$, so $\lambda<\lambda^{*}$.
@@
Let $c^{1}$ and $c^{2}$ represent two arbitrary peripheral-center edges or cycles on $G$. Note that $c^{1}$ is $c^{2}$ if all necessary centers covering $\mathcal{P}$ under $\lambda$ are in the same cycle or out-of-cycle edge. Let $T^{1}$ and $T^{2}$ represent the two subtrees of $T$ that respectively consist of $c^{1}$ and $c^{2}$ (i.e., their representations on $T$). Initially, we set $T^{1}=T^{2}=T$, and then assign a flag $f$ to every node $u$ of $T^{1}$ (resp., $T^{2}$) so that $u$’s $f$ is true iff the hanging subtrees of $u$ in $T/T^{1}$ (resp., $T/T^{2}$) must contain center(s) under $\lambda$. The flag $f$ of every node in $T^{1}$ (resp., $T^{2}$) is initialized as false.
@@
We shall now present how to find a peripheral-center edge or cycle node on $T^{1}$ and $T^{2}$ in turn. This routine needs the supports of Lemmas 5 and 6 whose proofs are delayed to Subsection 3.1.
@@
Compute the centroid $c$ on $T^{1}$ which is a node of $T^{1}$ whose hanging subtrees each is of size no more than half of $T^{1}$ and can be found in $O(|T^{1}|)$ time [11, 12]. Then, if $c$ is a hinge or $G$-node, by applying Lemma 5 to $c$, we determine whether a center must be placed at $G_{c}$ and if not, which hanging subtree of $c$ contains a peripheral-center edge or cycle. If only $c$ is returned, which means that $G_{c}$ must be a center, then we set $T^{1}$ as $c$.
@@
Generally, Lemma 5 returns one or two nodes that are adjacent to $c$ in $T^{1}$. In the former case where only one adjacent node of $c$ is returned, the corresponding hanging subtree of $c$ on $T$ contains all necessary centers. Hence, we find in $T^{1}$ $c$’s hanging subtree including this node and reset $T^{1}$ as this subtree. The latter case means that the two corresponding hanging subtrees of $c$ on $T$ each contains a peripheral-center edge or cycle. So, we find in $O(|T^{1}|)$ time the two hanging subtrees of $c$ in $T^{1}$ respectively containing the two obtained nodes. Set $T^{1}$ as the one of them that excludes any node with a true flag. Last we set $c$’s flag $f$ in $T^{1}$ as true.
@@
On the other hand, $c$ is a cycle node. Instead, we apply Lemma 6 to decide whether $G_{c}$ is a peripheral-center cycle and which hanging subtree of $c$’s $H$-subtree $T^{H}_{c}$ in $T^{1}$ contains a peripheral-center edge or cycle. As described in the proof for Lemma 6, if $c$ is obtained, which means that $G_{c}$ is a peripheral-center cycle, then we let $T^{1}$ be $c$. If one of its adjacent hinge nodes is returned, which can be known in $O(|T^{1}|)$ time, then a center must be placed at that hinge; so we set $T^{1}$ as that hinge node.
@@
Otherwise, we obtain two nodes that are respectively in two split subtrees of $T^{H}_{c}$ and adjacent to $c$’s hinge node(s), or obtain only one such node. For the former situation, we find in linear time a hanging subtree of $T^{H}_{c}$ in $T^{1}$ that includes one of the two nodes but excludes any node with a true flag $f$; next we set $T^{1}$ as this subtree and set the flag $f$ as true of $c$’s hinge node in $T^{1}$. For the latter case, we find the hanging subtree of $T^{H}_{c}$ in $T^{1}$ that contains the only obtained node and reset $T^{1}$ as this subtree.
@@
We recursively apply the above procedure to $T^{1}$ to find a peripheral-center edge or cycle until $T^{1}$ consists of only one node or one edge. Clearly, it takes $O(\log mn)$ recursive steps, so the running time is $O(mn\log mn)$.
@@
    $T^{1}$ is a hinge node or $G$-node. A center must be placed at vertex $G_{T^{1}}$ on $G$, which is represented by $T^{1}$. We can decide the feasibility of $\lambda$ in $O(mn\log mn)$ time as follows. Set the weight of every uncertain point that can be covered by $G_{T^{1}}$ under $\lambda$ as zero. Then we apply the one-center algorithm [9] to solve the one-center problem of $G$ w.r.t. $\mathcal{P}$ in $O(mn\log mn)$ time. If the obtained objective value is not greater than $\lambda$, then $G_{T^{1}}$ and the obtained center can cover $\mathcal{P}$ completely under $\lambda$, which means that $\lambda$ is feasible. Otherwise, $\lambda$ is not feasible.
@@
    $T^{1}$ is a cycle node. By the definition, we first determine which hanging subtree of $T^{1}$’s $H$-subtree contains a peripheral-center edge or cycle. If none, then all necessary centers can be placed on $G_{T^{1}}$. Hence, Lemma 8 presented in Subsection 3.1 is called to decide the feasibility of $\lambda$ in $O(m^{2}n^{2})$ time. Otherwise, we let $T^{2}$ be that hanging subtree of $T^{1}$’s $H$-subtree in $T$ and set the flag $f$ as true for the hinge node of $T^{2}$ that connects $T^{1}$ and $T^{2}$ in $T$, which can be carried out in $O(|T|)$ time.
@@
    $T^{1}$ is an edge. If it is between a hinge and a cycle node, clearly, this cycle node represents a peripheral-center cycle. We set $T^{1}$ as this cycle node and then handle this case in the above way in $O(m^{2}n^{2})$ time. Otherwise, $T^{1}$ is an out-of-cycle edge on $G$. We call Lemma 7 in Subsection 3.1 to decide the feasibility of $\lambda$ in $O(mn\log mn)$ time.
In general, $T^{1}$ is a cycle node representing a peripheral-center cycle, i.e., $c^{1}$, and $T^{2}$ is a subtree of $T$ which contains no vertices in $T^{1}$ and where a leaf connects it and $T^{1}$ on $T$ and is of flag $f$ being true. $G_{T^{2}}$ contains all other necessary centers under $\lambda$.
Proceed with finding a peripheral-center edge or cycle on $T^{2}$ recursively in the above way. After $O(\log mn)$ recursive steps, $T^{2}$ shrinks down to a node or an edge on $T$. If it is an out-of-cycle edge, we call Lemma 7 to decide the feasibility of $\lambda$ in $O(mn\log mn)$ time. Otherwise, $T^{2}$ contains exactly one cycle node, which is not same as $T^{1}$; $c^{2}$ is the cycle on $G$ that $T^{2}$ represents. Subsequently, Lemma 9 in Subsection 3.1 is applied to $T^{1}$ and $T^{2}$, which decides in $O(mn^{2})$ time the feasibility of $\lambda$.